Solve the equation by completing the square. The equation has real number solutions.
x^2+4x-2=0

Found 2 solutions by stanbon, jorel1380:
Answer by stanbon(75887)

Solve the equation by completing the square. The equation has real number solutions.
x^2+4x-2=0
-------
x^2 + 4x + 4 = 2+4
-------
(x+2)^2 = 6
------
x+2 = sqrt(6) or x+2 = -sqrt(6)
------
x = -2+sqrt(6) or x = -2-sqrt(6)
-----------
Cheers,
Stan H.
-----------

Answer by jorel1380(3719)

If x²+4x-2=0, then:
x²+4x=2
x²+4x+4=6
(x+2)²=6
x+2=+/-√6
x=(+/-√6)-2. ☺☺☺☺